1) phosphorus pentachloride is produced by the reaction between phosphine and chlorine gas. What mass of chlorine gas would be required to produce 0.015kg of phosphorus pentachloride
1
Expert's answer
2019-04-02T05:58:36-0400

The balanced equation for considered reaction can be written as follows

"PH_3 (g) + 4 Cl_2 (g) \rightarrow PCl_5 (g) + 3 HCl (l)"

"\frac{m(Cl_2)}{M(Cl_2) \cdot n(Cl_2)} = \frac{m(PCl_5)}{M(PCl_5) \cdot n(PCl_5)}"

"m(Cl_2) = \frac{m(PCl_5) \cdot M(Cl_2) \cdot n(Cl_2)}{M(PCl_5) \cdot n(PCl_5)}"

"M(Cl_2) = 2 \cdot 35.453 = 70.906 \; g/mol"

"M(PCl_5) = 1 \cdot 30.974 + 5 \cdot 35.453 = 208.239 \; g/mol"

"m(Cl_2) = \frac{0.015 \cdot 70.906 \cdot 4}{208.239 \cdot 1} \approx 0.0204 \; kg = 20.4 \; g"


Answer: 20.4 g.